# What Matters in Transformers? Not All Attention is Needed

### v3

@@ 5 Conclusion
In this work, we systematically revisit transformer architectures through the lens of module dropping, exploring three different structures: Blocks, MLP layers, and Attention layers. Our findings reveal that attention layers exhibit surprisingly high redundancy and can be removed in large proportions without compromising performance. Additionally, we further propose Joint Layer Drop, which further promotes the dropping ratios and performance. Our study empirically demonstrates the potential to make transformer models more compact and efficient, providing valuable insights for future network design within the NLP community.
@@ 6 Limitations
Despite the progress we have made, there are still limitations in our work. While our discussed dropping techniques demonstrate improved efficiency on the models we examined, a potential future work is a further exploration to verify their application across a broader range of models, e.g., vision transformers and vision-language models. Additionally, our work mainly focuses on post-training dropping and does not involve retraining, which has the potential to recover or even enhance performance further. Our work provides valuable insights for network architecture design, and we believe that incorporating further training will contribute to superior performance in more compact architectures.
@@ Appendix A Implementation Details
We utilize Llama-2 Touvron et al. (2023) and Mistral Jiang et al. (2023) as the default models, given their competitive performance and wide usage. We also evaluated the completely open-source model MAP-Neo Zhang et al. (2024) to explore the redundancy variations in the modules of the entire pre-training phase. Additionally, we experimented with the newly released Llama-3 to verify the effectiveness of model dropping on the latest models.
@@
For the calibration dataset, we used the validation set of C4 dataset Raffel et al. (2019), with 256 samples and an input sequence length of 2,048, following the setup in Sun et al. (2023). To evaluate the performance of the model, we report the results of the following tasks: BoolQ Clark et al. (2019), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), ARC-C Clark et al. (2018), HellaSwagZellers et al. (2019), MMLU Hendrycks et al. (2021), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). Please refer to Table 4 for detailed information. The evaluation code is based on EleutherAI LM Evaluation Harness Gao et al. (2023).
Table 4: Parameter settings for each task. “Norm” refers to the normalization performed with respect to the length of the input.
@@ Appendix B Additional Experimental Results
Figure 9: Accuracy Curves on GSM8k, where Mistral-7B-Base represents mistralai/Mistral-7B-v0.1, Llama-2-7B-Base represents meta-llama/Llama-2-7b-hf, Mistral-7B-Math represents meta-math/MetaMath-Mistral-7B, and Llama-2-7B-Math represents meta-math/MetaMath-7B-V1.0.
@@
To evaluate Attention Drop on more complex technical tasks, we evaluated Llama-2-7B and and Mistral-7B, and two corresponding instruction fine-tuned models, meta-math/MetaMath-7B-V1.0 and meta-math/MetaMath-Mistral-7B Yu et al. (2023). The results in Figure 9 indicate that, except for Llama-2-7B-Math, which is meta-math/MetaMath-7B-V1.0, all the models do not experience significant performance degradation when dropping fewer than 8 Attention layers. We speculate that this is because Llama-2-7B-Math is initialized with Llama-2-7B and undergoes instruction fine-tuning to improve its mathematical ability. Instead, Llama-2-7B-Base exhibits poor performance in mathematics, and the ability obtained solely through fine-tuning appears to be superficial. Therefore, when dropping Attention layers, Llama-2-7B-Math’s ability rapidly deteriorates.
@@
Simple One-Shot Drop is Effective.
Iterative and One-Shot are the most common settings for model compression, where the former prunes step by step and the latter prunes the model once. Iterative pruning may causes more accurate measurements but tend to be more time-consuming. In Figure 10, we empirically compare Iterative Dropping and One-Shot Dropping. For Iterative Dropping, we consider dropping one by one iteratively.
However, based on Figure 10, Iterative Dropping achieves performance that is merely comparable to One-Shot Dropping, without offering any significant enhancement. Given its simplicity and efficiency, One-Shot Dropping emerges as the superior choice.
@@
Given that quantization simplifies data types and enhances efficiency in memory usage and inference speed, we integrate module dropping with quantization to verify whether the quantized models can maintain the performance achieved by Attention Drop. Specifically, we use the mainstream AWQ algorithm Lin et al. (2024) for 4-bit quantization, following its default settings, which involve using 128 samples from the Pile dataset Gao et al. (2020) as the calibration dataset.
@@
Based on Figure 11 and 12, similar to Attention Drop, the importance scores for MLP Drop and Block Drop also demonstrate robustness to calibration datasets, regardless of changes in data type or sample size. Although minor differences exist in the scores, the resulting dropped modules are nearly identical.
@@
Figure 11: The Impact of Sample Quantity on the Importance Scores of Block and MLP.
@@
Figure 12: The Impact of Dataset Type on the Importance Scores of Block and MLP.

### v4

@@ 5 Conclusion
In this work, we systematically revisit transformer architectures through the lens of module dropping, exploring structured redundancy across three different modules: Blocks, MLP layers, and Attention layers. Our findings reveal that attention layers exhibit surprisingly high redundancy and can be removed in substantial proportions without compromising performance. Additionally, we introduce Joint Layer Drop, which enhances both dropping ratios and overall performance. Our study empirically demonstrates the potential to make transformer models more compact and efficient, offering valuable insights for future network design within the AI community.
@@ 6 Limitations
Despite the progress made, there are still some limitations in our work. While we have identified redundancy in attention layers within mainstream large language models, future research could explore whether similar findings hold across a broader range of models, including vision transformers and vision-language models. Additionally, while our study primarily focuses on training-free module dropping, post-training has the potential to recover or even promote performance. We believe that incorporating post-training training techniques could lead to better performance.
@@ Appendix A Implementation Details
We utilize Llama-2 Touvron et al. (2023) and Mistral Jiang et al. (2023) as our default models due to their competitive performance and wide usage. To explore the redundancy dynamics in modules throughout the pretraining phase, we also considered the fully open-source model MAP-Neo Zhang et al. (2024). Additionally, we experimented with the newly released Llama-3 to assess the effectiveness of model dropping on the latest models.
@@
For the calibration dataset, we used the validation set of the C4 dataset Raffel et al. (2019), with 256 samples and an input sequence length of 2,048, following the setup in Sun et al. (2023). To evaluate the performance of the model, we report the results of the following tasks: BoolQ Clark et al. (2019), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), ARC-C Clark et al. (2018), HellaSwagZellers et al. (2019), MMLU Hendrycks et al. (2021), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). Please refer to Table 4 for detailed information. The evaluation code is based on EleutherAI LM Evaluation Harness Gao et al. (2023).
Table 4: Experimental settings for evaluation tasks. “Norm” refers to the normalization performed with respect to the length of the input.
@@ Appendix B Additional Experimental Results
Figure 9: Accuracy Curves on GSM8k.
@@
To evaluate Attention Drop on more complex technical tasks, we evaluated Llama-2-7B and and Mistral-7B, and two corresponding instruction fine-tuned models, meta-math/MetaMath-7B-V1.0 and meta-math/MetaMath-Mistral-7B Yu et al. (2023). The results in Figure 9 indicate that, except for Llama-2-7B-Math, all the models do not experience significant performance degradation when dropping fewer than 8 Attention layers. We speculate that this is because Llama-2-7B-Math is initialized with Llama-2-7B and undergoes instruction fine-tuning to improve its mathematical ability. Instead, Llama-2-7B-Base exhibits poor performance in mathematics, and the ability obtained solely through fine-tuning appears to be superficial. Therefore, when dropping Attention layers, Llama-2-7B-Math’s ability rapidly deteriorates.
@@
Given that quantization enhances efficiency and is orthogonal with pruning, we further integrate Attention Drop with quantization. Specifically, we use the mainstream AWQ algorithm Lin et al. (2024) for 4-bit quantization, following its default settings, which involve using 128 samples from the Pile dataset Gao et al. (2020) as the calibration dataset.
@@
Effectiveness of One-Shot Dropping.
Iterative and One-Shot are two common settings for model compression. The Iterative method prunes the model step by step, allowing for more precise measurements but often requiring more time. In contrast, the One-Shot method prunes the model in a single step. Figure 10 presents an empirical comparison between Iterative Dropping and One-Shot Dropping. For Iterative Dropping, we drop modules one by one in a sequential manner.
However, as shown in Figure 10, Iterative Dropping only achieves performance that is merely comparable to One-Shot Dropping, without offering any significant enhancement. Given its simplicity and efficiency, One-Shot Dropping emerges as the superior choice.
@@
As shown in Figure 11 and 12, similar to Attention Drop, the importance scores for MLP Drop and Block Drop also demonstrate robustness to calibration datasets, regardless of changes in data type or sample size. Although minor differences exist in the scores, the resulting dropped modules are nearly identical.
@@
Figure 11: Impact of Sample Quantity on the Importance Scores of Block and MLP.
@@
Figure 12: Impact of Dataset Type on the Importance Scores of Block and MLP.